How do you rewrite an equation that has both x and y values? For example, r=34k+56t. How do you solve this so that k is a function of r, and then k as a function of t.
Answer by Alan3354(69443)

How do you rewrite an equation that has both x and y values?
2x + 3y = 7 is one.
-------------------
For example, r=34k+56t.
Not a good example, it doesn't have x or y.
--------------
How do you solve this so that k is a function of r, and then k as a function of t.
You can solve for k, but it will be a function of both r and t, since they're both in the given equation.
r=34k+56t
34k = r-56t
k = (r-56t)/34